• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 10:24:38 on May 09, 2011

    Entry number 358709

    This entry is a followup to: 358705

    Followups:

    keyword=L1A delays of both two HRSs have not been change since Mar4.

    I looked at my note and also the trigger files in atrig@adaql5:/trigger/,
    and I found that actually I did not change the delay values of L1A when
    switching from SRC to x>2 experiment. The last modification date of
    trigger_left.map and trigger_right.map was Mar. 4th. I can not remember
    whether at that time we thought we did not have to change the delays or
    we changed the L1A delay at the bigbite weldment. Alexandre and I set up
    two scopes for two HRSs to check the trigger timing. You might want to
    bring up those two screen through the website and you should see whether
    the retiming function works properly or not. Alexandre knows the IP
    address for both. 
    
    I don't know why there is a 60ns shift. My suggestion is still not to see
    the RAW TDC values but check the real timing you care, before you take an
    access to check connections. 
    


    A copy of this log entry has been emailed to: vasulk,camsonne